OpenStudy (anonymous):

please answer this for me. if i have a 54 in my class and my final exam is worth 425 points what do i need to get on the exam to have a 75 in the class?

OpenStudy (dan815):

you have a 54 what?

OpenStudy (anonymous):

How many points is the total class?

OpenStudy (anonymous):

like a 54.85 i think and i have no idea 100 i assume

OpenStudy (anonymous):

You can't figure this out without knowing the amount of points total in the course... sorry

OpenStudy (anonymous):

what does that mean exactly? point total in the course? and can you like take a guess?

OpenStudy (anonymous):

I can't because I don't know how many points are in the course. You've earned 54.85% of the points possible. I need to know how many points you'ld have if you had earned 100% of the points possible.

OpenStudy (anonymous):

oh i would have a 100 if i had 100 precent of the point possible

OpenStudy (dan815):

what grade are you in

OpenStudy (anonymous):

bout to be senior

OpenStudy (anonymous):

Okay out then in total you have 525 points possible. In order to get a 75 in the class you must get 525(.75) points. Solve that

OpenStudy (anonymous):

Then subtract the points you already have and thats the total of points you need inorder to get 75% in the class

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

Okay The test is worth 425 points and the rest of the class was worth 100 points. The total would be 525. 75% of 525=393.75 that means you need 393.75 points in order to get 75% in the class. You've already earned 54.85 points so 393.75-54.85=338.9 in order to get 75% in the class you need 338.9 points on the test.

OpenStudy (anonymous):

better?

OpenStudy (anonymous):

ok so if there are 40 questions on the final how many do i have to get right to get the 75?

OpenStudy (anonymous):

425/40=10.625 Each question is worth 10.625 points 338.9/10.625= 31.9 Round that up and you get that you need at least 32 questions right.

OpenStudy (anonymous):

ok thank you very much

OpenStudy (anonymous):

No prob c: